# Some stability results for the fractional differential equations with two delays

### v1

@@ 5.1 Illustrative Examples for Case 2
Consider the FDDE $D^{\alpha}x(t)=-\gamma x(t)+kx(t-1)-ke^{-\gamma\tau}x(t-1-\tau)$.
@@
Take $\alpha=0.9,\;k=2.5,\;\gamma=8.5,\;\tau=0.6$. Since $\gamma>2k>0$, Theorem 5.1(a)(i) guarantees delay-independent stability. Simulations confirm convergence (see Figure (16).
Figure 16: $\alpha=0.9,\;k=2.5,\;\gamma=8.5,\;\tau=0.6$
Example 5.1.2 Now, consider $\alpha=0.7,\;k=-3,\;\gamma=7.5,\;\tau=0.4$. Here $\gamma>-2k>0$. By Theorem 5.1(a)(ii), the system is asymptotically stable for all $\tau$. Numerical solutions confirm this (cf. Figure (17).
Figure 17: $\alpha=0.7,\;k=-3,\;\gamma=7.5,\;\tau=0.4$
Example 5.1.3 For $\alpha=0.8,\;k=-4,\;\gamma=-7,\;\tau=0.08$. Since $k<0,\;\gamma<0$, Theorem 5.1(b) predicts delay-independent instability. The simulations show divergence.
Figure 18: $\alpha=0.8,\;k=-4,\;\gamma=-7,\;\tau=0.08$
@@
Example 5.1.4 For $\alpha=0.8,\;k=6,\;\gamma=-0.4,\;\tau=1.6$. Since $k>0,\;\gamma<0$, according to Theorem 5.2, there exists $k_{*}$ such that equilibrium is unstable for all $k>k_{*}.$ Now, $k_{*}=5.675449$, so, k>kk>k*. Hence, the equilibrium is unstable.
Figure 19: $\alpha=0.8,\;k=6,\;\gamma=-0.4,\;\tau=1.6$
@@ 5.2 Illustrative Examples for a Nonlinear Function $g(x)$
$D^{\alpha}x(t)=-\gamma x(t)+k\sin\!\bigl(x(t-1)\bigr)-e^{-\gamma\tau_{2}}\,k\sin\!\bigl(x(t-1-\tau)\bigr),\quad 0<\alpha\leq 1.$ (13)
@@
$D^{\alpha}x(t)=-\gamma x(t)+kx(t-1)-ke^{-\gamma\tau}x(t-1-\tau).$
@@
Example 5.2.1. Let $\alpha=0.35,\;k=2.4,\;\gamma=6,\;\tau=0.55$. Then, equation (13) becomes
$D^{0.35}x(t)=-6x(t)+2.4\sin(t-1)-2.4e^{-3.3}\sin(t-1.55).$
The solutions are stable (cf. Fig. LABEL:3a), in agreement with Theorem 5.1, since here $\gamma>2k>0$.
Figure 20: $\alpha=0.35,\;k=2.4,\;\gamma=6,\;\tau=0.55$.
Example 5.2.2. Now take $\alpha=0.65,\;k=-0.8,\;\gamma=2,\;\tau=1.6$. In this case,
$D^{0.65}x(t)=-2x(t)-0.8\sin(t-1)+0.8e^{-3.2}\sin(t-2.6).$
@@
Figure 21: $\alpha=0.65,\;k=-0.8,\;\gamma=2,\;\tau=1.6$.
Example 5.2.3. Finally, consider $\alpha=0.88,\;k=-3.5,\;\gamma=-5.7,\;\tau=0.85$. Equation (13) becomes
$D^{0.88}x(t)=5.7x(t)-3.5\sin(t-1)+3.5e^{5.7\cdot 0.85}\sin(t-1.85).$
@@
Figure 22: $\alpha=0.88,\;k=-3.5,\;\gamma=-5.7,\;\tau=0.85$.
@@ 6 Conclusion
This paper investigated the stability of a fractional-order delay differential equation with a delay-dependent coefficient and two discrete delays. By linearizing the system and analyzing the characteristic equation, we derived both delay-independent and delay-dependent stability conditions. We studied two cases: $\tau_{1}=0$ and $\tau_{1}=1$ in detail. The numerical simulations and stability diagrams supported the theoretical findings.

### v2

@@ 5.1 Illustrative Examples for Case 2
Consider the FDDE $D^{\alpha}x(t)=-\gamma x(t)+kx(t-\tau_{1})-ke^{-\gamma\tau_{2}}x(t-\tau_{1}-\tau_{2})$.
@@
Take $\alpha=0.33,\;k=2.6,\;\gamma=6.4,\;\tau_{1}=0.35,\;\tau_{2}=0.45$. Since $\gamma>2k>0$, Theorem 5.1(a)(i) guarantees delay-independent stability. Simulations confirm convergence (see Figure (16(a))).
(a) $\alpha=0.33,\;k=2.6,\;\gamma=6.4,\;\tau=0.35,\;\tau_{2}=0.45$
(b) $\alpha=0.8,\;k=6.4,\;\gamma=15,\;\tau=0.85,\;\tau_{2}=1.45$
Figure 16: $\gamma>2k>0$
Example 5.1.2
Take $\alpha=0.8,\;k=6.4,\;\gamma=15,\;\tau_{1}=0.85,\;\tau_{2}=1.45$. Again, $\gamma>2k>0$. So, equilibrium is stable for these values of parameters (cf. Figure(16(b))).
Example 5.1.3
Now, consider $\alpha=0.56,\;k=-4.9,\;\gamma=12,\;\tau_{1}=1.37,\;\tau_{2}=1.49$. Here $\gamma>-2k>0$. By Theorem 5.1(a)(ii), the system is asymptotically stable for all $\tau$. Numerical solutions confirm this (cf. Figure (17(a))).
(a) $\alpha=0.56,\;k=-4.9,\;\gamma=12,\;\tau_{1}=1.37,\;\tau_{2}=1.49$
(b) $\alpha=0.78,\;k=-5.2,\;\gamma=13.4,\;\tau_{1}=1.25,\;\tau_{2}=0.64$
Figure 17: $\gamma>-2k>0$
Example 5.1.4
Take $\alpha=0.78,\;k=-5.2,\;\gamma=13.4,\;\tau_{1}=1.25,\;\tau_{2}=0.64$. Again, $\gamma>-2k>0$. So, equilibrium is stable (cf. Figure(17(b))).
Example 5.1.5
For $\alpha=0.65,\;k=-2.35,\;\gamma=-4.56,\;\tau_{1}=0.82,\;\tau_{2}=0.41$. Since $k<0,\;\gamma<0$, Theorem 5.1(b) predicts delay-independent instability. The simulations show divergence (see Figure (18(a))).
(a) $\alpha=0.65,\;k=-2.35,\;\gamma=-4.56,\;\tau_{1}=0.82,\;\tau_{2}=0.41$
(b) $\alpha=0.89,\;k=-3.31,\;\gamma=-8.41,\;\tau_{1}=0.42,\;\tau_{2}=0.62$
Figure 18: $k<0,\;\gamma<0$
Example 5.1.6
Now, take $\alpha=0.89,\;k=-3.31,\;\gamma=-8.41,\;\tau_{1}=0.42,\;\tau_{2}=0.62$. Here, again $(k,\gamma)$ lies in the third quadrant. So, according to Theorem 5.1(b) equilibrium lies in the unstable region (cf. Figure (18(b)).
@@
Example 5.1.7
For $\alpha=0.43,\;k=14,\;\gamma=-2.32,\;\tau_{1}=1.28,\;\tau_{2}=2.35$. Since $k>0,\;\gamma<0$, according to Theorem 5.2, there exists $k_{*}$ such that equilibrium is unstable for all $k>k_{*}.$ Now, $k_{*}=11.3536$, so, k>kk>k*. Hence, the equilibrium is unstable.
Figure 19: $\alpha=0.43,\;k=14,\;\gamma=-2.32,\;\tau_{1}=1.28,\;\tau_{2}=2.35$
@@ 5.2 Illustrative Examples for a Nonlinear Function $g(x)$
$D^{\alpha}x(t)=-\gamma x(t)+k\sin\!\bigl(x(t-\tau_{1})\bigr)-e^{-\gamma\tau_{2}}\,k\sin\!\bigl(x(t-\tau_{1}-\tau_{2})\bigr),\quad 0<\alpha\leq 1.$ (14)
@@
$D^{\alpha}x(t)=-\gamma x(t)+kx(t-\tau_{1})-ke^{-\gamma\tau_{2}}x(t-\tau_{1}-\tau_{2}).$
@@
Example 5.2.1. Let $\alpha=0.35,\;k=2.4,\;\gamma=6,\;\tau_{1}=0.82,\;\tau_{2}=1.49$. Then, equation (14) becomes
$D^{0.35}x(t)=-6x(t)+2.4\sin(t-0.48)-2.4e^{-8.94}\sin(t-2.32).$
The solutions are stable (cf. Fig. 20), in agreement with Theorem 5.1, since here $\gamma>2k>0$.
Figure 20: $\alpha=0.35,\;k=2.4,\;\gamma=6,\;\tau_{1}=0.48,\;\tau_{2}=1.29$.
Example 5.2.2. Now take $\alpha=0.65,\;k=-0.8,\;\gamma=2,\;\tau_{1}=0.79,\;\tau_{2}=0.38$. In this case,
$D^{0.65}x(t)=-2x(t)-0.8\sin(t-0.79)+0.8e^{-0.76}\sin(t-1.17).$
@@
Figure 21: $\alpha=0.65,\;k=-0.8,\;\gamma=2,\;\tau_{1}=0.79,\;\tau_{2}=0.38$.
Example 5.2.3. Finally, consider $\alpha=0.88,\;k=-3.5,\;\gamma=-5.7,\;\tau_{1}=1.62,\;\tau_{2}=0.85$. Equation (14) becomes
$D^{0.88}x(t)=5.7x(t)-3.5\sin(t-1.62)+3.5e^{4.845}\sin(t-2.47).$
@@
Figure 22: $\alpha=0.88,\;k=-3.5,\;\gamma=-5.7,\;\tau_{1}=1.62,\;\tau_{2}=0.85$.
Acknowledgment
Pragati Dutta thanks the University of Hyderabad for the non-net fellowship (UH/DSW/F&/Non-NET/2023/165).
@@ 6 Conclusion
This paper investigated the stability of a fractional-order delay differential equation with a delay-dependent coefficient and two discrete delays. By linearizing the system and analyzing the characteristic equation, we derived both delay-independent and delay-dependent stability conditions. The numerical simulations and stability diagrams supported the theoretical findings.